Please help with this question, it is an online course so I don't have an ISBN number.
A sample of 25 concession stand purchases at the October 22 matinee of Bride of Chucky showed a mean purchase of $5.29 with a standard deviation of $3.02. For the October 26 evening showing of the same movie, for a samle of 25 purchases the mean was $5.12 with a standard deviation of $2.14 the means appear to be very close but not the variances. At alpha=.05, is there a difference in variances? Show an illustration of the decision rule. thank you!

Answer by stanbon(48568)

A sample of 25 concession stand purchases at the October 22 matinee of Bride of Chucky showed a mean purchase of $5.29 with a standard deviation of $3.02. For the October 26 evening showing of the same movie, for a sample of 25 purchases the mean was $5.12 with a standard deviation of $2.14 the means appear to be very close but not the variances. At alpha=.05, is there a difference in variances? Show an illustration of the decision rule. thank you!
-----------------------
Ho: variance(matinee) = variance(evening)
Ha: the variances are not equal
----------------
Matinee variance = 3.02^2 = 9.1204
Evening variance = 2.14^2 = 4.5796
----------
Critical Value for alpha=5%,2-tail F test, df=24 for numerator; df=24 for denominator:
F =0.44 and 2.27
-----------
Test statistic: F = 4.5796/9.1204 = 0.5021...
=========================
Since 0.44 < 0.5021 < 2.27, Fail to Reject Ho.
Conclusion: Test statistic is in the "Fail to Reject" interval.
The test provides evidence that the two variances are statistically
equal.
---------------------------
Cheers,
Stan H.